OpenStudy (anonymous):

The formula for the area of a trapezoid is A=1/2h(a+b), where h represents the height and a and b represent the length of the bases. Solve for h.

OpenStudy (anonymous):

\(\large A=\frac{1}{2}h(a+b) \) \(\large 2A=h(a+b) \) multiply 2 to both sides; \(\large \frac{2A}{a+b}=h \) divide both sides by a+b

OpenStudy (anonymous):

where did you get 2 from?

OpenStudy (anonymous):

cause the original function is A=1/2h(a+b), then u multiply both sides with 2...then u get 2A=1/2*2*h(a+b)